Can you please help me answer this question.
x/3=4/x + 1

Answer by stanbon(75887)

Can you please help me answer this question.
x/3=4/x + 1
---
Multiply thru by 3x to get:
x^2 = 12 + 3x
x^2 - 3x -12 = 0
Use the quadratic formula:
x = [3 +- sqrt(9-4*-12)]/2
x = [3 +- sqrt(57)]/2
x = (3/2) +- (1/2)sqrt(57)
=============================
Cheers,
Stan H.